# A Generative Framework for Probabilistic, Spatiotemporally Coherent Downscaling of Climate Simulation

### v3

@@ Abstract
Figure 1: Probabilistic pipeline for spatiotemporal downscaling of multiple variables. This work introduces a probabilistic downscaling pipeline that jointly estimates spatio-temporally consistent weather dynamics from ESM simulations on multiple variables. The framework is built around a score-based diffusion model and can be understood as a combination of four modules, which can each be adjusted independently of the others. This schematic outlines the framework. Only one exemplary variable is shown for visual clarity. A: A diffusion model is trained on reanalysis data. For a pre-defined set of variables, it learns to reproduce the spatial and temporal patterns from the reanalysis data it is trained on. Note that ESM simulations are not part of the training process. B: From any ESM (e.g. CMIP6), select and pre-process an ensemble run (e.g. MPI-HR) for downscaling. A bias-correction step that partially mitigates distributional deviations between simulated climate and reanalysis data can be applied to the climate input as a pre-processing step. C: Formulate a known or assumed functional relationship between climate and weather dynamics. This step imposes a constraint onto the generative model such that its samples adhere to the encoded prior knowledge. D: The model generates weather trajectories that preserve the statistics of the coarse climate input. During the generative process, the trained score model (A) is conditioned to fulfill the constraint imposed by explicitly relating the conditioning information (B) to the generated samples through the observation model (C).
@@
Numerical simulations based on the Navier-Stokes equations, discretized over time and space, are fundamental to understanding weather patterns, climate variability, and climate change. State-of-the-art numerical weather prediction (NWP) models, which primarily focus on atmospheric processes, can accurately resolve small-scale dynamics within the Earth system, providing fine-scale spatial and temporal weather patterns at resolutions on the order of kilometers [1]. However, the substantial computational resources required for these models render them impractical for simulating the extended time scales associated with climatic changes. In contrast, Earth System Models (ESMs), such as those included in the CMIP6 project [2], incorporate a broader range of processes—including atmospheric, oceanic, and biogeochemical interactions—while operating on coarser spatial scales. Typical grid resolutions for ESMs are approximately $1^{\circ}$, equivalent to around 100 km. This coarse resolution limits the ability of ESMs to fully capture small-scale processes, requiring parameterizations to represent unresolved dynamics as functions of resolved variables. These parameterizations reduce the local detail in ESM outputs, introducing a scale mismatch between modeled outputs and the fine-scale processes that influence local climate. Consequently, ESM data cannot be directly employed to evaluate changes at the local scale, restricting their utility for localized decision-making. To address this limitation, novel approaches are needed to infer local dynamics from coarse-grained models while preserving their global-scale information. Machine learning (ML), particularly techniques derived from image generation, offers a promising avenue for overcoming these challenges. Statistical downscaling methods powered by ML enhance the spatial and temporal resolution of ESM outputs, providing detailed local information and addressing the limitations of traditional modeling approaches. Despite their promise, applying ML to climate modeling presents a significant challenge. ESMs are designed to generate accurate multi-decadal summary statistics, which inherently differ from historical observational datasets. This misalignment complicates the training of ML models, which typically require well-aligned datasets to optimize the model by minimizing the difference between simulations and observations. Generative models, a subset of ML techniques, offer a potential solution. Unlike other ML approaches, generative models learn the underlying data distribution, enabling them to generate plausible sample trajectories. Recent methodological advances in generative modeling [3, 4, 5], in particular a diffusion model designed to emulate physically plausible dynamics that are conditioned on known or assumed relationships between observed and predicted quantities [6], allows bridging the gap between coarse-resolution ESM outputs and fine-scale local processes, thereby enhancing the utility of ESMs for localized and regional analyses, even in the context of future climate projections.
We leverage and extend the score-based diffusion generative model of Rozet and Louppe [6] (figure 1) to (simultaneously) increase the spatial and temporal resolution of multivariate coarse climate model output. We show that this approach enables high-fidelity realizations of local weather patterns. The model is first trained on high-resolution reanalysis data to capture the statistical and physical properties of local weather dynamics, after which it is conditioned on climate simulations for spatial and temporal downscaling. This framework requires an explicitly defined functional relationship between coarse-grained climate inputs and high-resolution weather outputs, thereby integrating prior knowledge into the modeling process. By separating the training of weather dynamics from task-specific inference, the model becomes highly flexible and versatile. This enables it to be applied to a range of tasks, including downscaling, which can be performed in a zero-shot manner without requiring retraining.
In our experiments, we demonstrate that this model generates weather sequences that are physically plausible, spatially and temporally consistent, and aligned with climate projections. The model effectively recovers local weather dynamics, including extreme events such as cyclones. Furthermore, its generative nature allows it to provide samples from the posterior distribution, enabling the generation of multiple plausible weather trajectories. This ability is crucial for assessing the variability and uncertainty, providing a comprehensive understanding of future weather scenarios.
@@ Results
Figure 2: Comparison of value distributions: reanalysis data, coarse input, predictions. For a time range of 49 hours, this plot shows that the 1-hourly local predictions, which the model predicted from the coarse 6-hourly input, resemble the reanalysis data closely, in distribution. Top row: Kernel-density estimations for value distributions of reanalysis data (green) and 30 predictions (black). The prediction model was conditioned on coarse inputs (purple). The predicted samples align with the reanalysis data distribution, which is fully covered by the predictive uncertainty. Bottom row: The probability integral transform (PIT) assesses the uncertainty calibration of the model: for each variable separately (eh) and overall (i). A PIT distribution that resembles a standard uniform distribution indicates that the reanalysis data and predictions likely come from the same distribution.
@@
We first evaluate the methodological framework used for downscaling in an on-model setting. We construct an experimental setup with artificially coarsened reanalysis data that surrogates the ESM simulations. This setting simulates a direct pairing between the climate and reanalysis data. It thus allows for evaluating the model’s predictive capabilities and uncertainty quantification by comparing it to a ground truth. This setup is described in the Methods section.
From the coarse input data, the model predicts weather trajectories that match the reanalysis data in distribution when aggregated across space and time. Figure 2 ad show that the estimated densities of the predictions cover those of the reanalysis data in the predicted uncertainty. In particular, the probabilistic model introduces no systematic biases, like distribution shift, a tendency towards over- or under-predicting values, or mismatch in the tail regions. Instead, each sampled prediction captures the spread of the data distribution. To assess the calibration of the predictive uncertainty, we compute the probability integral transform (PIT) for all values aggregated and for each variable separately. We find that both wind-speed components (figure 2 g,h) are well calibrated. The predictions for mean sea-level pressure (figure 2 e) are slightly too defensive, over-predicting extreme values. For surface temperature (figure 2 f), the opposite is the case: the predicted distribution is slightly too narrow and under-predicts tail events. Overall (figure 2 i), the model is well calibrated.
@@
Recovering local dynamics from coarse information
@@
Recovering complex local dynamics from coarse information is a crucial requirement we impose on the downscaled predictions. This experiment shows that our model accurately predicts spatial and temporal variations in weather trajectories at two distinct locations, which share a single spatial observation. To this end, we investigate two locations with distinct weather dynamics that both lie in a region that corresponds to one node on the coarse observation grid (figure 3 e). To ensure the model can recover meaningful local information from coarse information, we compare the temporal evolution of the downscaled trajectories at both locations in figure 3. We find that the predictions’ time series align with the reanalysis data throughout the considered temporal period and across all variables. As an illustrative example, the zonal wind speeds (figure 3 c), show substantial variability between the two locations—both in their values and their temporal structure. The model captures these variations in its predictions, highlighting its ability to recover local-scale variability lost in the coarse-resolution observations.
@@
Figure 4: Predicting high-resolution dynamics during a cyclone as an extreme event. This plot shows meridional wind trajectories of three randomly selected model predictions (top three rows), reanalysis data (fourth row), and coarse conditioning information (bottom row) during a cyclone ("Friederike", January 2018). The sign of the wind speed value defines its direction: negative values go southward, and positive values go northward. Time progresses from left to right hourly, starting 2018 January 18 at 02:00 PM and ending the same day at 08:00 PM. Between the first (2:00 PM) and the last (8:00 PM) visualized time point, no information is provided to the model. The top three rows show how different samples fill in the missing spatial and temporal information. Each individual generated trajectory aligns visually with the coarse input, while the variation among the samples captures the uncertainty associated with the inference problem. Notably, the model does not introduce implausible “jumps” from one observation to the next but interpolates with spatially and temporally consistent dynamics.
Spatiotemporally consistent weather trajectories
@@
Statistical downscaling is particularly challenging during extreme events since they are rare and thus lie in the tails of the training distribution while potentially involving highly complex dynamics. However, extreme events are particularly interesting, as they often have the most significant societal and environmental impacts. To assess the model performance in reproducing extreme events, we choose the time range in which cyclone "Friederike" approached central Europe (including the modeled spatial region) from the west around January 18, 2018. Aside from the time period, the experimental setup remains unaltered from the above sections. This setting is especially challenging during a storm event since local spatial and temporal dynamics are hidden from the model and have to be inferred based on very scarce information. However, the sub-sampled observations are enough to guide the model in accurately predicting the cyclone. Figure 4 shows the meridional wind trajectories during the event for three randomly selected model predictions, the reanalysis data, and the coarse conditioning information. While the samples have significant variation, all three predictions are consistent in space and time and plausibly explain the provided coarse observations. Furthermore, the radially-averaged power spectral density of the predictions closely matches the reanalysis data (supplementary figure 2(a)), showing the physical plausibility of the predictions.
Local climate information through downscaling
@@
Figure 5: Comparison of value distributions: ESM, de-biased ESM, reanalysis data, predictions. This plot shows two things: firstly, the bias-correction reduces the distributional shift between ESM output and reanalysis data. Secondly, the model can generate high-resolution weather trajectories that introduce complex local weather dynamics while preserving the value distribution of the ESM simulations. Comparing the top row and the middle row shows that the distribution mismatch between both ESM ensembles and the reanalysis data (ad) is mitigated through the bias-correction step (eh). The bottom row shows that the distribution of the downscaled climate trajectories aligns with the coarse model input. The value histograms for the de-biased ESM data in the bottom row are copied over from the middle row. The considered time range is the year 2014; precisely 1-hourly steps starting January 01 at 06:00 AM and ending December 31 at 06:00 AM. Two distinct CMIP6 ensembles (MPI and HadGEM) are visualized together, which allows a comparison of the distributional mismatch between a) the respective climate simulations and b) the climate projections and the reanalysis data. For each ESM, eight downscaled samples are shown. The model does not introduce systematic biases, the downscaled ESM distributions are not shifted or skewed unexpectedly, and the tails align with their coarse counterparts.
@@
Figure 6: Local climate information through downscaling. The plot shows spatiotemporally downscaled predictions for the MPI-HR ensemble that is part of the CMIP6 simulations. Time progresses from left to right in three-hourly steps, starting 2014 August 24 at 06:00 AM and ending the same day at 03:00 PM. The four variables, mean sea-level pressure (a), surface temperature (b), and zonal (c) and meridional (d) wind-speed components are downscaled jointly by the model. The top three rows show the progression of three randomly selected samples. The bottom row shows the corresponding conditioning information from the coarse bias-corrected (BC) ESM simulations. Where the observation is blank, the model interpolates in time, without any conditioning information.
Through a series of experiments, we demonstrated that the model can infer small-scale dynamics from artificially coarsened data such that the predictions are spatially, temporally and physically plausible while preserving the underlying value distribution. Building on these findings, we now move to downscaling climate outputs, using 6-hourly CMIP6 ESM simulations as conditioning information. As in the previous experiments, the four variables considered are downscaled spatially (by a factor of $16\times 16$) and temporally (by a factor of $6$) to align with the resolution of the reanalysis data.
While this setup is similar to the previous experiments, differing only in the type of conditioning data, the problem is more challenging. Climate simulations and reanalysis data are not aligned; hence, we do not have access to ground truth (paired high- and low-resolution) data. Furthermore, raw climate model outputs typically contain systematic biases when compared to observational data. And ESM models are, of course, themselves subject to simulation error. As illustrated in the top row in figure 5, these biases result in systematic deviations between ESM outputs and reanalysis data distributions. We mitigate climate model biases using a per-variable quantile-mapping approach [7]. Figure 5 (ad $\mapsto$ eh) shows that the bias-correction step reduces the distributional shift between the climate and weather distribution.
Reducing the discrepancies between climate and weather distributions is not the task of the generative downscaling model but is instead addressed by a bias correction applied to the conditioning input. However, bias correction cannot resolve fundamental simulation errors in climate models [8]. Climate models are inherently imperfect representations of the climate system, with inaccuracies in parameterizations, simplified process representations, and uncertainties regarding the system’s initial state [9]. These variations lead to significant discrepancies between models; choosing a model, therefore, becomes a critical factor [10]. Even in the hypothetical case of perfect models (i.e., no epistemic uncertainty), forecasts are not deterministic due to the chaotic nature of the atmosphere. The internal variability of the climate system leads to differences between projected and observed climate [11]. This implies that climate simulations represent only one possible realization of the system with substantial uncertainty remaining. To illustrate this, figure 5 shows two model ensembles from different climate models: the MPI-HR and the HadGEM model. This illustrates that, even in terms of their aggregated-value distribution, climate model outputs can differ substantially from other ensembles and the reanalysis data.
@@
Figure 6 presents a qualitative assessment of the spatial and temporal progression of downscaled ESM simulations. The model output reproduces spatial patterns, such as the Alps, and temporal patterns, like the day-night cycle, while introducing local weather dynamics consistent with observations and propagating in a physically plausible manner. Together, figure 5 and figure 6 show that the model can generate high-resolution weather trajectories that introduce complex local dynamics while preserving the value distribution of the respective ESM simulations. Climate information at high spatial and temporal resolution is extremely valuable for understanding future climate changes and variability, especially for sectors like renewable energy. Our model predicts downscaled ESM simulations that enable more accurate local wind-energy forecasts, illustrated in the supplementary material (figures A.7 and A.8).

### v4

@@ Abstract
Introduction
@@
Numerical weather and climate simulations based on discretized solutions of the Navier-Stokes equations are fundamental to understanding large-scale weather patterns, climate variability, and climate change. State-of-the-art numerical weather prediction (NWP) models, which primarily focus on atmospheric processes, can accurately resolve small-scale dynamics within the Earth system, providing fine-scale spatial and temporal weather patterns at resolutions on the order of kilometers [1]. However, the substantial computational resources required for these models render them impractical for simulating the extended time scales of multiple years and decades necessary to assess climatic changes. Moreover, even with substantial computational investment, global high-resolution models can still exhibit systematic biases and may fail to accurately reproduce observed climatic trends [2]. In contrast, earth system models (ESMs), such as those included in the CMIP6 project [3], incorporate a broader range of processes—including atmospheric, oceanic, and biogeochemical interactions—while operating on coarser spatial scales. Typical grid resolutions for ESMs are approximately $1^{\circ}$, equivalent to around 100 km. This coarse resolution limits the ability of ESMs to fully capture small-scale processes. Key processes necessary to assess regional impact, for example, on wind turbines—such as local wind turbulence—occur at spatial and temporal scales that are too fine to be explicitly resolved in ESMs. Consequently, ESM data cannot be directly employed to evaluate changes at fine spatial scales, limiting their utility for localized impact assessment and decision-making.
Downscaling aims to provide regional climate information by estimating small-scale processes from coarse simulations of global models. Existing approaches to bridge this scale gap can be categorized into dynamical and statistical downscaling [4]. Dynamical downscaling employs high-resolution regional climate models (RCMs) that are nested within coarser global climate models (GCMs) to provide detailed projections for specific regions [5, 6]. Biases from the driving global models can be inherited by the RCMs, potentially limiting the accuracy of the downscaled results [7]. Additionally, the high computational cost of running RCMs restricts their use primarily to regional studies. Statistical downscaling methods use regression or weather generators [8], and, more recently, emulators based on machine learning (ML) techniques. In statistical downscaling, a functional or statistical relationship between large-scale climate variables (from GCMs) and local observations is established. Based on this relationship, statistical methods infer the local information from the coarse simulation, with ML-based approaches aiming to capture the mapping from coarse to fine-scale climate by learning the relationships from data [9, 10, 11, 12]. These models are usually computationally less expensive. However, not explicitly encoding physical laws can make them less robust in regions of low data density, risking physical inconsistencies.
Figure 1: Probabilistic pipeline for spatiotemporal downscaling of multiple variables. This work introduces a probabilistic downscaling framework that jointly predicts fine-scale and spatiotemporally consistent time series for multiple variables from coarse ESM simulations. This schematic outlines the framework. Only one exemplary variable is shown for visual clarity. A: A score model is trained on sequences of reanalysis data. This is the centerpiece of the diffusion model that learns to reproduce the fine-scale spatial and temporal patterns. Note that ESM simulations are not part of the training process. B: From any ESM (e.g. CMIP6), select and pre-process an ensemble run (e.g. MPI-HR) for downscaling. A bias-correction step that mitigates distributional deviations between climate output and reanalysis data can be applied as a pre-processing step. C: This part establishes a relationship between coarse climate simulations and the (fine-scale) output space of the model. The observation model defines how the observed quantity ($Y_{\mathrm{ESM}}$) is generated—or observed—from the inferred quantity ($X_{\mathrm{reanalysis}}$). The observation model is key to impose a constraint onto the generative model such that its samples adhere to the established relationship. D: The model generates time series that preserve the statistics of the coarse climate input. During the generative process, the trained score model (A) is conditioned, i.e., the predictions are informed by the conditioning information (B), such that they adhere to the relationship established by the observation model (C).
One challenge that comes with downscaling arises from the fact that numerical models are inherently imperfect representations of the climate system. While ESMs are designed to generate accurate multi-decadal summary statistics, locally, the simulations differ from historical observational datasets. Climate models contain inaccuracies in parameterization, simplified process representations, uncertainties in the initial state of the system [13]. Even in the hypothetical case of perfect models, i.e., without epistemic uncertainty, forecasts are not deterministic due to the chaotic nature of the atmosphere. These variations lead to substantial discrepancies between models. Choosing a model, therefore, becomes a critical factor [14, 15]. Aside from other climate model biases [16], internal variability of the climate system leads to differences between projected and observed climate [17]. Individual climate simulations thus represent only one possible realization of the system with substantial uncertainty remaining. Due to this un-pairedness of ESM outputs and observational data, using supervised ML approaches, which rely on consistent simulation–observation pairs, remains challenging [18, 19]. Generative models have recently emerged as a promising solution. This model class is characterized by learning a representation of the training data distribution that allows the generation of novel samples. As self-supervised learning techniques, these models circumvent the need for data–label pairs by working solely on the target (output) distribution. Additionally, through the variability among generated samples, generative models provide structured uncertainty, which the ill-posed nature of most inference problems entails. In particular, diffusion models (DMs) [20, 21, 22, 23] have demonstrated superior performance over earlier approaches such as variational auto-encoders (VAEs) [24], generative adversarial networks (GANs) [25], and normalizing flows [26], particularly for structured data and image generation tasks [27, 28, 29].
In this work, we build on the score-based data assimilation (SDA) framework by Rozet and Louppe [23, 30] to phrase downscaling as a Bayesian-inference problem with a generative prior model. By construction, our model performs joint spatial and temporal downscaling on multiple variables, introducing stochasticity solely between samples. The predictions are thus coherent across the entire state space, avoiding sampling-induced inconsistencies between time steps and between interrelated atmospheric variables. Furthermore, the model training is separated from the task-specific inference (Fig. 1 I & II), which makes the model flexible with respect to its input and the statistical relationship between input and output—allowing, for instance, downscaling climate simulations of varying spatiotemporal resolutions without retraining the model. The components of the algorithmic pipeline are outlined in Figure˜1:
  • A
    The score model, which is the centerpiece of the generative diffusion model, is trained on high-resolution reanalysis data. The score-based diffusion model provides a statistical representation of local weather dynamics. This "prior" model gives access to samples from the distribution $p(X_{\mathrm{reanalysis}})$, which represents our prior concept of the output space. Here, the output space includes the spatial region, the target resolution, and learned dynamics patterns for a set of selected atmospheric variables.
  • B
    The coarse ESM input $Y_{\mathrm{ESM}}$ is pre-processed; in particular, a bias-correction procedure can align the simulation with the reanalysis data in terms of its value distribution.
  • C
    The "observation model" $p(Y_{\mathrm{ESM}}\mid X_{\mathrm{reanalysis}})$ establishes the functional or statistical relationship between coarse climate output and fine reanalysis data. This assumes the ESM output to be a perfect prognosis [31, 32], based on which local climate is estimated. In the context of inverse problems, the observation model is often referred to as the "forward model", as it models how the partially observed quantity ($Y_{\mathrm{ESM}}$) arises from the latent quantity of interest ($X_{\mathrm{reanalysis}}$)—usually by removing information.
  • D
    Accordingly, the inverse problem aims to predict the (much harder, underspecified) opposite direction: estimating the "posterior" $p(X_{\mathrm{reanalysis}}\mid Y_{\mathrm{ESM}})$ requires adding information to the incomplete observations by conditioning the prior. In our case, the trained score model is coupled with the observation model to generate samples from the posterior distribution.
Perspectively, the algorithmic framework is versatile and flexible enough to be considered beyond downscaling. In the spirit of foundation models [33, 34, 35, 36], the trained generative model can be combined with other observation models in a similar zero-shot manner to solve various inference tasks on the target output space. Unlike approaches that require training the model directly on the conditioning information, the presented framework allows the formulation of explicit—and varying—functional or statistical relationships between observations and predictions.
On a series of experiments, we demonstrate that the proposed model generates coherent time series of regional climate that are aligned with coarse input. The model predicts local weather dynamics, including extreme events such as winter storms. Sampling from the posterior distribution enables the generation of multiple weather trajectories crucial for assessing the internal variability and uncertainty of the downscaling problem, providing a comprehensive understanding of future weather scenarios. We include a simple quantile-mapping procedure as a pre-processing step to the ESM simulations. Other than that, the ESM simulations are taken as a perfect predictor and the mismatch to the observational data is not accounted for. We evaluate our downscaling framework on two different GCMs from CMIP6 in order to show that different mismatches between reanalysis and the respective ESM distributions remain. We then demonstrate that our model maintains the coarse, global properties of the respective ESM input while inferring local fine-scale weather patterns.
@@ Results
The presented downscaling pipeline assumes a statistical relationship between a coarse numerical model prediction and a fine-scale reanalysis product [31, 32]. We begin by evaluating the methodological framework in an artificial setup, in which coarsened reanalysis data serve as the perfect prediction [37] and surrogate the ESM simulations. The coarse data is obtained from spatial area averages and by selecting a subset of the time steps. Concretely, the observation model establishes the following relationship between the downscaled predictions $X$ and the coarse input $Y$:
$Y^{(i,j)}_{t}=\frac{1}{\lvert\overline{i}\rvert\cdot\lvert\overline{j}\rvert}\cdot\sum_{(i^{\prime},j^{\prime})\in\overline{i}\times\overline{j}}X^{(i^{\prime},j^{\prime})}_{t}$, (1)
where $(i,j)$ is a single point on the coarse spatial grid, which represents an area of multiple points $\overline{i}\times\overline{j}$ on the fine grid. For example, $\overline{i}$$\overline{j}$ can be a local neighborhood around ii/$j$ in the longitude/latitude dimension. The observation model in Equation˜1 assumes that the coarse information is provided as a snapshot at time points $t$. In our experiments, one coarse-grid point $(i,j)$ encompasses a $16\times 16$-area on the fine grid, i.e., $\lvert\overline{i}\rvert\cdot\lvert\overline{j}\rvert=16^{2}=256$. Furthermore, the temporal resolution of the coarse grid is six hours, whereas the fine output time grid is resolved hourly. This setting enables a direct pairing between climate output and reanalysis data, which is not given in reality [37], allowing for evaluating the model’s predictive performance and uncertainty quantification by comparing it to a ground truth. Having established the validity of the downscaling method on this artificial setup, we will then proceed to downscale climate output, assuming two different ESM simulations from the CMIP6 project to be the perfect predictors.
@@
Figure 2: Comparison of value distributions: reanalysis data, coarse input, predictions. For a time range of 49 hours, this plot shows that the 1-hourly local predictions, which the model predicted from the coarse 6-hourly input, resemble the reanalysis data closely in distribution. Top row: Kernel-density estimations for value distributions of reanalysis data (green) and 30 predictions (black). The prediction model was conditioned on coarse inputs (purple). The predicted samples align with the reanalysis data distribution, which is fully covered by the predictive uncertainty. Bottom row: The probability integral transform (PIT) demonstrates the uncertainty calibration of the model: for each variable separately (eh) and overall (i). A PIT distribution that resembles a standard uniform distribution indicates that the reanalysis data and predictions likely come from the same distribution.
We first evaluate the basic capability of the model to adhere to the established statistical relationship between coarse and fine predictor (cf. Equation˜1) and to preserve the coarse-data value distribution in its predictions. Aggregating the values over temporal and spatial domains, we find that—per variable—the estimated densities of the predictions each align with the reanalysis data. The uncertainty induced by the sample spread covers the reanalysis distribution both near the modes and in regions of low and high quantiles. In particular, the probabilistic model introduces no systematic biases, like distribution shift, a tendency towards over- or under-predicting values, or mismatch in the tail regions. Instead, each sampled prediction captures the spread of the data distribution. Figure˜2 ad visualizes these findings using density estimations of the respective value distributions.
To assess the calibration of the predictive uncertainty, we compute the probability integral transform (PIT) [38] for all values aggregated and for each variable separately. If the resulting distribution (Figure˜2 ei) resembles a standard uniform distribution on the interval $[0,1]$, it is likely that the samples and the reanalysis data come from the same underlying distribution. We find that both wind-speed components (Figure˜2 g,h) are well calibrated. The PIT reveals that the predictions for mean sea-level pressure (Figure˜2 e) are slightly underconfident, over-predicting extreme values. For surface temperature (Figure˜2 f), the opposite is the case: the predicted distribution is slightly too narrow and under-predicts tail events. Taken the joint distribution of variables together (Figure˜2 i), the model is well calibrated.
Predicting local dynamics from coarse information
@@
We require our downscaling model to augment the scarce information contained in the coarse input by adding nontrivial, local weather patterns and predicting complex temporal and spatial dynamics. As described above, in our experimental setup, a single scalar measurement informs a six-hour window of $16\times 16$ fine-grid locations, requiring the model to perform a mapping from a single node to $6\times 16\times 16=1\,536$ nodes. Hence, we need to assess whether the model sensibly incorporates prior knowledge, which it learned from data, in order to evaluate the plausibility of the added, generated information.
It is likely that—due to, for instance, varying environmental conditions—the dynamical patterns at distinct locations on the reanalysis grid differ substantially from each other, whereas the coarse-grid aggregation occludes these local variations, motivating the downscaling problem in the first place. The present input-output-paired setup allows us to investigate the true local variations that are lost through aggregation by comparing the reanalysis data at two distant locations within the $16$-by-$16$ area that is encompassed by a single coarse-grid location. Accordingly, we can compare these ground-truth variations with our model predictions to assess whether the model infers local dynamics that align with the reanalysis data.
@@
In Figure˜3, we demonstrate that our model accurately predicts spatial and temporal variations in weather trajectories at two distinct locations, which share a single spatial observation. For this experiment, we selected two distant fine-grid locations near the Alps, for which we can expect substantial local variations (see purple circle and cross in Figure˜3 e). We visualize the time series for the reanalysis data at both locations (solid and dotted green line) and the corresponding downscaled model predictions (solid and dotted black lines) alongside the coarsened input (purple circle and cross), which share a single value at every 6-hourly step. The visualization allows to identify the local variations that are occluded in the spatiotemporal aggregation by comparing the coarse observations (purple) to the fine ground-truth time series. Our model accurately predicts the ground-truth dynamics at both locations from the coarse information. Notably, the distinct features in temporal structure (e.g. smoothness and amplitude) of the predicted time series at both locations align with those found in the corresponding reanalysis data. As an illustrative example, the zonal wind speeds (Figure˜3 c) show substantial variability between the two locations—both in their values and their temporal structure. The model captures these variations in its predictions, highlighting its ability to recover local-scale variability lost in the coarse-resolution observations. This is a promising indicator that the downscaling model has learned an accurate representation of local spatial and temporal patterns from the training data, which it blends into the coarse conditioning information to estimate coarsely informed local variations.
Spatiotemporally consistent weather trajectories: studying a winter storm
@@
Figure 4: Predicting high-resolution dynamics during a cyclone as an extreme event. This plot shows meridional wind trajectories of three randomly selected model predictions (top three rows), spatiotemporal interpolation (fourth row), reanalysis data (fifth row), and coarse conditioning information (bottom row) during a cyclone ("Friederike", January 2018). The sign of the wind speed value defines its direction: negative values go southward, and positive values go northward. Time progresses from left to right hourly, starting 2018 January 18 at 02:$0$ PM and ending the same day at 08:$0$ PM. Between the first (2:$0$ PM) and the last (8:$0$ PM) visualized time point, no information is provided to the model. The top three rows show how different samples add missing spatial and temporal information, while the interpolation (fourth row) can only spread out the existing, coarse information. Each individual generated trajectory aligns visually with the coarse input, while the variation among the samples captures the uncertainty associated with the inference problem. Notably, the model does not introduce implausible “jumps” from one time step to the next but interpolates with spatially and temporally consistent dynamics.
@@
Statistical downscaling becomes particularly challenging during extreme events, which are rare and fall into the tails of the training distribution. These events often involve highly complex dynamics. However, extreme events are particularly interesting, as they often have the most significant societal and environmental impacts. As Figure˜2 already demonstrated, the downscaling model is capable of accurately predicting high quantiles. Here, we supplement the evaluation of the aggregated value distributions with a qualitative assessment of the spatiotemporal structure of the downscaled predictions during a winter storm. To this end, we consider the time range in which cyclone "Friederike" approached central Europe (including the modeled spatial region) from the west around January 18, 2018. Aside from the time period, the experimental setup remains unaltered from the above sections. Figure˜4 exemplary shows the meridional wind speeds during the event. The visualization compares three randomly selected model predictions, a spatiotemporal interpolation of the coarse input, the ground-truth reanalysis data, and the coarse input. The model predictions are coherent in space and time and add different local variations to the coarse observations. A comparison to the spatiotemporal interpolation (Figure˜4, fourth row) highlights the capability of the downscaling model to predict nontrivial fine-scale patterns. Supplementary Figure 9 visualizes anomalies—differences between the interpolation and a) downscaled predictions and b) reanalysis data. This visualization reveals the spatiotemporal patterns, which are lost by aggregating the reanalysis data, and allows a comparison to the disaggregated fine-scale structure predicted by our downscaling model. Additionally, Supplementary Figure 2 demonstrates a close match of power spectral densities between reanalysis data and model predictions, providing a more quantitative argument.
Downscaling ESM simulations
@@
Figure 5: Comparison of value distributions: ESM, de-biased ESM, reanalysis data, predictions. This plot visualizes aggregated value distributions of reanalysis data (green), ESM simulations (two different models: MPI in purple, HadGEM in yellow), and downscaled predictions (black). The top row compares the raw, uncorrected ESM simulations with the reanalysis data to visualize the biases in the respective climate distributions. The middle row shows the bias-adjusted ESM simulations alongside the same reanalysis data to visualize the effect of the quantile-mapping procedure. The bottom row compares the value distribution of the bias-adjusted ESM simulations (same as middle row) with their downscaled counterpart (black). The eight solid/dotted lines correspond to downscaled predictions of the MPI/HadGEM model. Two things are demonstrated: Comparing the top row and the middle row shows that the distribution mismatch between both ESM ensembles and the reanalysis data (ad) is mitigated through the bias-correction step (eh). There is some mismatch remaining, likely due to the short evaluation period. Secondly, the bottom row shows that the distribution of the downscaled climate trajectories aligns with the coarse model input. The considered time range is the year 2014. The model predicts 1-hourly steps starting January 01 at 06:00 AM and ending December 31 at 06:00 AM based on the corresponding 6-hourly coarse input. Visualizing two distinct CMIP6 ensembles (MPI and HadGEM) allows a comparison of the distributional mismatch a) between the respective climate outputs (purple vs. yellow) and b) between climate outputs and reanalysis data.
The above experiments evaluated the model in a setting that ensures that the large-scale predictor perfectly matches the statistics of the fine-scale reanalysis data. We proceed to downscaling climate-model outputs, using 6-hourly CMIP6 ESM simulations as conditioning information. As in the previous experiments, the four variables considered are downscaled spatially (by a factor of $16\times 16$) and temporally (by a factor of $6$) to align with the resolution of the reanalysis data. The same statistical relationship (Equation˜1) between coarse input and downscaled output is assumed in this experiment. Downscaling climate-model outputs, however, presents a greater challenge because there is no direct pairing between coarse- and fine-scale climate data, which renders the comparison of our downscaling predictions to a ground truth impossible. The goal of this experiment is to predict local patterns on a fine spatiotemporal grid, assuming two distinct perfect predictors given by two realizations from the distribution of coarse climate simulations. We evaluate whether our downscaling model—while adhering to the statistical relationship imposed between each prediction and different coarse ESM simulations (Figure˜5)—is capable of simultaneously predicting local climate (Figure˜6). Notably, climate model biases are not encoded directly and therefore not addressed by the downscaling model. As motivated by Volosciuk et al. [39], downscaling can be separated into, firstly, mitigating climate model biases and, secondly, bridging the gap between coarse and fine grids. We adopt this perspective, thereby focusing almost exclusively on the latter. As a pre-processing step, we apply a per-variable quantile-mapping bias-correction [40] to the ESM outputs.
@@
Figure 6: Local climate information through downscaling. The plot shows spatiotemporally downscaled predictions for the MPI-HR ensemble that is part of the CMIP6 simulations. Time progresses from left to right in three-hourly steps, starting on August 24, 2014, at 06:00 AM and ending the same day at 03:$0$ PM. The four variables, mean sea-level pressure (a), surface temperature (b), zonal (c) and meridional (d) wind-speed components, are downscaled jointly by the model. The top three rows show the progression of three randomly selected samples. The bottom row shows the corresponding conditioning information from the coarse bias-corrected (BC) ESM simulations. Where the observation is blank, the model interpolates in time, without any conditioning information.
Figure˜5 ad shows two model ensembles of the four variables from different climate models: the MPI-HR (purple) and the HadGEM model (yellow) for the year 2014 and, in comparison, the reanalysis data distribution (green curve). Comparing the distributions of the two ESMs, it becomes evident that, even in terms of their aggregated-value distribution, climate model outputs can differ significantly from each other and, moreover, each individual ESM differs substantially from the reanalysis data. Figure˜5 eh visualizes the effect of the quantile-mapping bias adjustment, which adjusts the marginal distributions of each variable in the ESM simulations to better match those of the reanalysis data. As a result, the bias-corrected ESM simulations more closely align with the reanalysis data (green curve). A comparison of the uncorrected and bias-corrected distributions (Figure˜5 ad vs. eh) reveals that some residual differences remain. Figure˜5 il visualizes the distribution of eight downscaled predictions (black) for each of the two climate models. The downscaled predictions for MPI-HR (solid lines) and HadGEM (dotted lines) closely match the value distributions of their respective bias-corrected ESM simulations. Consequently, the downscaling model preserves the statistical properties of the coarse ESM input: the downscaled distributions are neither shifted nor skewed relative to their bias-corrected coarse counterparts, and the alignment extends to the tails of the distributions. As a result, any temporal changes or variations in the distribution shape present in the coarse input are faithfully retained in the downscaled output.
@@
Figure˜6 presents a qualitative assessment of the spatial and temporal progression of downscaled ESM simulations. The model output reproduces spatial patterns, which are, for example, consistent with geographical features, such as the Alps, and temporal patterns, like the day-night cycle, while introducing local weather dynamics consistent with the observations. Together, Figure˜5 and Figure˜6 show that the model can generate high-resolution weather trajectories that introduce complex local dynamics while preserving the value distribution of the respective ESM simulations.